If two bulbs of wattages 25 W and 100 W respectively , each rated at 220 V are connected in series with the supply of 440 V, which bulb will get fused?
1
Expert's answer
2012-07-05T09:45:20-0400
Let's find out the resistances of each bulb from the data given:

R1 = U1^2 / P1 = 220^2 / 25 = 1936 Ohm

R2 = U2^2 / P2 = 220^2 / 100 = 484 Ohm
Let's use P = I^2 * R for each bulb and find out if it is greater than its power rating. The one with higher
power rating will get fused.


P1 = (440 / (1936 + 484))^2 * 1936 = 64 W

P2 = (440 / (1936 + 484))^2 * 484 = 16 W.

Answer: first bulb will get fused (64W > 25W).